# Scalable and consistent embedding of probability measures into Hilbert spaces via measure quantization

### v1

@@ 3.2.3 Clustering performances
$\mathrm{WCSS}(l,\nu_{K})\leq 3\mathrm{WCSS}(l,\mu)+\frac{6N}{N_{l}}\varepsilon_{K}.$ (19)
@@
$\mathrm{BCSS}(l_{1},l_{2},\nu_{K})\geq\frac{1}{3}\mathrm{BCSS}(l_{1},l_{2},\mu)-\Bigl{(}\frac{N}{N_{l_{1}}}+\frac{N}{N_{l_{2}}}\Bigr{)}\varepsilon_{K},$ (20)
@@
3.3 Machine learning from the quantized measures
We consider here the embeddings of measures into a Hilbert space using either LOT or KME in Section 2. We focus on comparing the performances of machine learning methods after such embeddings of the input measures $(\mu^{(i)})_{i=1}^{N}$ and their quantized versions $(\nu^{(i)}_{K})_{i=1}^{N}$. To this end, we compare the Gram matrices of the pairwise inner-products between the set of embedded measures. Indeed, these matrices play a crucial role in various machine learning tasks [22] such as PCA or Linear Discriminant Analysis (LDA), that rely on the diagonalization of the covariance operator of data in a Hilbert space, which is equivalent to diagonalizing the Gram matrix of inner-products as recalled in Appendix B.
In the following, for a given embedding $\phi$ of probability measures into a Hilbert space $\mathcal{H}$ equipped with the inner-product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$, we will denote $G^{\phi}_{\mu}$ and $G^{\phi}_{\nu_{K}}$ the $N\times N$ Gram matrices associated with $(\mu^{(i)})_{1\leq i\leq N}$ and $(\nu^{(i)}_{K})_{1\leq i\leq N}$ respectively, with entries
$(G^{\phi}_{\mu})_{ij}=\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}},\mbox{and}\quad(G^{\phi}_{\nu_{K}})_{ij}=\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}},$
and $\|\cdot\|_{F}$ the Frobenius matrix norm.
@@ 3.3 Machine learning from the quantized measures
$\frac{1}{N}\|G^{\mathrm{LOT}}_{\mu}-G^{\mathrm{LOT}}_{\nu_{K}}\|_{F}^{2}\leq C_{N,\mathcal{X},L}\sqrt{\varepsilon_{K}}$ (21)
@@
(ii) We note $G^{\mathrm{KME}}_{\mu}$ and $G^{\mathrm{KME}}_{\nu_{K}}$ the Gram matrices corresponding to the KME $\phi:\sigma\mapsto\int k(x,\cdot)\mathrm{d}\sigma(x)$ for a bounded kernel function $k$, then:
$\frac{1}{N}\|G^{\mathrm{KME}}_{\mu}-G^{\mathrm{KME}}_{\nu_{K}}\|_{F}^{2}\leq C_{N,k}\varepsilon_{K}$ (22)
where $C_{N,k}$ is a constant depending on $N$ and the kernel $k$.
@@ 4 Numerical experiments
In our experiments, we distinguish the two quantization approaches : $\tilde{K}$-LOT and $\tilde{K}$-KME refer to the LOT embedding and KME of $(\tilde{\nu}^{(i)}_{K})_{i=1}^{K}$ obtained from the quantization of each $\mu^{(i)}$, while $\overline{K}$-LOT and $\overline{K}$-KME refer to the LOT embedding and KME of $(\overline{\nu}^{(i)}_{K})_{i=1}^{K}$ obtained from the mean-measure quantization. Here, we aim to show that our method enables fast computation of machine learning tasks while preserving the main information of the measures.
@@ 4.1 Computational cost
For a discrete measure, we solve the quantization problem (8) using Lloyd’s algorithm [29] and an initialization based on $K$-means+++++ + [4]. The time complexity of the Lloyd’s algorithm being linear in the number of data points [21], it follows that for discrete measures $\mu_{1},\ldots,\mu_{N}$ supported on $m_{1},\ldots,m_{N}$ points respectively, the computational cost for constructing the quantized measures $(\nu^{(i)}_{K})_{i=1}^{K}$ by either optimal quantization of each input measure or mean-measure quantization is $O(Kd\sum_{i=1}^{N}m_{i})$. Nevertheless, as the support of $\overline{\mu}$ is very large in applications, the mean measure quantization can be done on $\frac{1}{N}\sum_{i=1}^{N}m_{i}$ points randomly sampled from $\overline{\mu}$, leading to optimal locations $\bar{x}_{1},\cdots,\bar{x}_{K}$ similar to those of the $K$-quantization of the entire support of $\overline{\mu}$. Then, the computation cost of mean-measure quantization becomes $O(Kd\frac{1}{N}\sum_{i=1}^{N}m_{i})$ and is advantageous over the optimal quantization of each input measure.
In order to compute the LOT embedding, we solve the discrete OT problem (2) between $m_{0}$ samples of $\rho$ and $m_{i}$ samples of $\mu^{(i)}$ using a standard OT solver [36, 15]. Then, as the optimal map $T_{\rho}^{\mu^{(i)}}$ in (5) might not exist, it is classical [13] to compute an approximation through barycentric projection of an optimal transport plan, solution of (2). The overall computational of $\overline{K}$-LOT when using $m_{0}$ samples from the reference measure $\rho$ is thus $O(Kd\frac{1}{N}\sum_{i=1}^{N}m_{i})+O(N(K+M)Km_{0}\log(K+m_{0}))$ which is significantly smaller than the one of LOT from the raw input measures that scales as $O(\sum_{i=1}^{N}(m_{i}+m_{0})m_{i}m_{0}\log(m_{i}+m_{0}))$.
@@ 4.2 Synthetic dataset : shifts and scalings of a reference measure
$\mu^{(i)}=(\Sigma_{i}^{1/2}\mathrm{id}+b_{i})_{\#}\rho,$ (23)
where $\Sigma_{i}\in\mathbb{R}^{d\times d}$ is a positive semi-definite matrix and $b_{i}\in\mathbb{R}^{d}$. We choose $X\sim\rho$ such that $X=R\frac{Z}{\|Z\|}$, with $R\sim\mathrm{Unif}([0,1])$ and $Z\sim\mathcal{N}(0,I_{d})$. In that case, we have an explicit formulation of the pairwise inner-products induced by LOT embedding and KME, see Proposition C.1 in Appendix C. We can therefore exactly compute the so-called true matrix $G_{\mu}^{\phi}$. We numerically sample the distributions $\mu^{(i)}$’s in the following way : for each $1\leq i\leq N$, we first sample $m_{i}$ points $(x_{j})_{1\leq j\leq m_{i}}$ from the measure $\rho$ by sampling $z_{j}\sim\mathcal{N}(0,I_{d})$ and $r_{j}\sim\mathrm{Unif}([0,1])$ and computing $x_{j}=r_{j}\frac{z_{j}}{\|z_{j}\|}$. This allows to sample points from the unit ball in $\mathbb{R}^{d}$. Samples from $\mu^{(i)}$ are then obtained by the pushforward operation in (23).
We first compare in Figure 2 the computational costs of both quantization methods and observe that the mean-measure quantization approach is faster. In Figure 3 (resp. Figure 4 in Appendix C), we visualize the projections on the first two components of PCA for $K$-LOT (resp. $K$-KME) of both quantization methods and compare them to the PCA computed from the true Gram matrices $G_{\mu}^{\phi}$ computed in Proposition C.1. We observe that even with a small value $K=32$, the PCA visualizations on quantized embedded measures mimic the ones on raw embedded input measures.
@@
Figure 3: Synthetic dataset on shifts and scalings. Projection of the data onto the first two components of PCA after $\overline{K}$-LOT (top) and $\tilde{K}$-LOT (bottom) and comparison to the LOT PCA (right) computed from the true Gram matrix (see Prop.C.1).
@@ 4.3 Flow cytometry dataset
Table 1: Flow cytometry dataset. Classification accuracies and execution times for LDA after 10-component PCA on ${K}$-LOT, ${K}$-KME with both quantization methods and KME with RFF.
$\overline{K}$-LOT/$\tilde{K}$-LOT $\overline{K}$-KME/$\tilde{K}$-KME KME with RFF
$K$ Accuracy (lab) Accuracy (type) Time (s) Accuracy (lab) Accuracy (type) Time (s) $s$ Accuracy (lab) Accuracy (type) Time (s)
16 100/100 85/81 23/103 100/100 83/69 15/96 16 73 44 4524
32 100/100 94/81 25/166 100/100 83/69 34/174 32 75 44 4701
64 100/100 94/81 30/281 100/100 85/69 105/358 64 83 52 5035
128 100/100 88/81 32/555 100/100 77/71 387/909 128 92 44 5676
@@
In this section, we use flow cytometry datasets provided in [43] and publicly available in Mendeley Data to illustrate the suitability of our method through a classification task. We have $N=108$ cytometry measures (or point cloud) which come from two different health care centers : Marburg and Dresden. In Marburg, the data consists of diagnostic samples of peripheral blood (pB), healthy bone marrow (BM), or leukemic bone marrow. The Dresden dataset consists of diagnostic samples of peripheral blood and healthy bone marrow. Two types of labels can be distinguished: data are differentiated either by the healthcare centers from which they were analyzed, or by their types (e.g., peripheral blood, healthy bone marrow, or leukemic bone marrow). Each measure contains from 100,000 to 1,000,000 points in dimension $d=10$, which prevents the use of classical OT, that is, without a quantization step. For $K$-LOT, we sample $m_{0}=1000$ points from the reference measure chosen as the uniform measure on $[0,1]^{d}$ and for $K$-KME, we use the Gaussian kernel with bandwidth parameter $\sigma=1$. The embedded measures live in a high-dimensional Hilbert space (e.g. with $K$-LOT and $K=128$, the ambient space is $\mathbb{R}^{1280}$). In order to keep the most relevant information, we perform a 10-components PCA on the embedded data with respect to the proper Hilbert space $\mathcal{H}$. Then, we train two classifiers (one for each type of label) on 75% of the data and test the results on the remaining data with LDA. We compare our method with random Fourier features (RFF) [38], which allow to efficiently map raw probability measures into the linear space $\mathbb{R}^{s}$, where $s$ is a parameters that has to be chosen. Note that RFF then approximate the KME. Accuracy scores are displayed in Table 1. The quantization methods achieve in particular perfect accuracy scores for predicting the health care center (denoted as LAB) in only a few minutes. The KME computed with RFF reaches lower accuracy scores and is clearly slower than both quantization methods. Comparing the quantization methods, we observe that the classifier performs better on mean-measure quantization than on the quantization of each measure whereas the latter is roughly ten times slower than the former.
@@
Additionally, we visualize the projections of the data on the first components of PCA in Figures 5 and 6 of the Appendix C. In these figures, it is clear that the representations stabilize when $K\geq 32$. This shows that our $K$-quantization step gives good results even for small values of $K$. We also observe that the mean-measure quantization and the quantization of each measure yield similar results.

### v3

@@ 4.3 Clustering performances
$\mathrm{WCSS}(l,\nu_{K})\leq 3\mathrm{WCSS}(l,\mu)+\frac{6N}{N_{l}}\varepsilon_{K}.$ (4.1)
@@
$\mathrm{BCSS}(l_{1},l_{2},\nu_{K})\geq\frac{1}{3}\mathrm{BCSS}(l_{1},l_{2},\mu)-\Bigl{(}\frac{N}{N_{l_{1}}}+\frac{N}{N_{l_{2}}}\Bigr{)}\varepsilon_{K},$
@@
4.4 Convergence of the Gram matrices of Hilbert space embeddings after quantization
We consider here the embedding of measures into a Hilbert space using either LOT or KME presented in Section 2. Given the embeddings of both the input measures $(\mu^{(i)})_{i=1}^{N}$ and their quantized approximations $(\nu^{(i)}_{K})_{i=1}^{N}$, we focus on comparing their performances on machine learning methods. To this end, we consider the Gram matrices (4.2) of the pairwise inner-products between the set of embedded measures . Indeed, these matrices play a crucial role in various machine learning tasks [25] such as PCA or Linear Discriminant Analysis (LDA), that rely on the diagonalization of the covariance operator of data in a Hilbert space, which is equivalent to diagonalizing the Gram matrix of inner-products as recalled in Appendix B.
In the following, for a given embedding $\phi:\mathcal{P}(\mathcal{X})\to\mathcal{H}$ of probability measures into a Hilbert space $\mathcal{H}$ equipped with the inner-product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$, we will denote $G^{\phi}_{\mu}$ and $G^{\phi}_{\nu_{K}}$ the $N\times N$ Gram matrices associated with $(\mu^{(i)})_{1\leq i\leq N}$ and $(\nu^{(i)}_{K})_{1\leq i\leq N}$ respectively, with entries
$(G^{\phi}_{\mu})_{ij}=\langle\phi(\mu^{(i)}),\phi(\mu^{(j)})\rangle_{\mathcal{H}},\mbox{and}\quad(G^{\phi}_{\nu_{K}})_{ij}=\langle\phi(\nu^{(i)}_{K}),\phi(\nu^{(j)}_{K})\rangle_{\mathcal{H}}.$ (4.2)
We also note $\|\cdot\|_{F}$ the Frobenius matrix norm.
@@ 4.4 Convergence of the Gram matrices of Hilbert space embeddings after quantization
$\frac{1}{N}\|G^{\mathrm{LOT}}_{\mu}-G^{\mathrm{LOT}}_{\nu_{K}}\|_{F}^{2}\leq C_{N,\mathcal{X},L}\sqrt{\varepsilon_{K}}$ (4.3)
@@
(ii) We note $G^{\mathrm{KME}}_{\mu}$ and $G^{\mathrm{KME}}_{\nu_{K}}$ the Gram matrices corresponding to the KME $\phi:\sigma\mapsto\int k(x,\cdot)\mathrm{d}\sigma(x)$ for a kernel function $k$. Assume that there exists a constant $C>0$ such that, for all $x,y\in\mathcal{X}$,
$k(x,x)+k(y,y)-2k(x,y)\leq C^{2}\|x-y\|^{2}.$
Assume also that $k$ is bounded by a constant $M_{k}<\infty$. Then,
$\frac{1}{N}\|G^{\mathrm{KME}}_{\mu}-G^{\mathrm{KME}}_{\nu_{K}}\|_{F}^{2}\leq C_{N,k}\varepsilon_{K}$ (4.4)
where $C_{N,k}$ is a constant depending on $N$, $M_{k}$ and $C$.
@@
4.5 Convergence of PCA in the Wasserstein sense
In this section, we fix a reference measure $\rho$ and we focus on the LOT embedding $\mu\mapsto T_{\rho}^{\mu}-id\in L^{2}(\rho)$. We note $T^{(i)}$ the Monge map between $\rho$ and $\mu^{(i)}$ and $T^{(i)}_{K}$ the Monge map between $\rho$ and $\nu^{(i)}_{K}$ obtained from $\mu^{(i)}$ with any of the two quantization methods presented in Section 3.1. To simplify notation, we note $\overline{T}=T-id$ for a Monge map $T$. In order to define the reconstuction error of PCA, we first define the covariance operators:
$\Sigma=\frac{1}{N}\sum\limits_{i=1}^{N}\overline{T}^{(i)}\otimes\overline{T}^{(i)},\qquad\mbox{and}\quad\Sigma_{K}=\frac{1}{N}\sum\limits_{i=1}^{N}\overline{T}^{(i)}_{K}\otimes\overline{T}^{(i)}_{K},$
where for all $f,g,h\in L^{2}(\rho),(f\otimes g)h=\langle f,h\rangle_{L^{2}(\rho)}g$. Both covariance operators admit spectral representations
$\Sigma=\sum\limits_{j\geq 0}\lambda_{j}P_{j},\qquad\mbox{ and}\quad\Sigma_{K}=\sum\limits_{j\geq 0}(\lambda_{K})_{j}(P_{K})_{j},$
where $(\lambda_{j})_{j\geq 0}$ and $((\lambda_{K})_{j})_{j\geq 0}$ are positive eigenvalues sorted in decreasing order, and the $P_{j}$’s and $(P_{K})_{j}$’s are rank one projectors. Borrowing notation from [43] that is focused in deriving non-asymptotic bounds on the reconstruction error of PCA in a Hilbert space, we write
$P^{\leq q}=\sum\limits_{j\leq q}P_{j},\qquad\mbox{and}\quad P_{K}^{\leq q}=\sum\limits_{j\leq q}(P_{K})_{j}$
the orthogonal projections onto the linear subspaces spanned by the first $q$ eigenvectors of $\Sigma$ and $\Sigma_{K}$. Let $\mathcal{P}_{q}=\{P:L^{2}(\rho)\rightarrow L^{2}(\rho)~{}|~{}P\text{~{}is orthogonal of rank}~{}q\}$. Given $P\in\mathcal{P}_{q}$, we define the reconstruction errors of PCA as:
$Q(P)=\frac{1}{N}\sum\limits_{i=1}^{N}\|\overline{T}^{(i)}-P\overline{T}^{(i)}\|_{L^{2}(\rho)}^{2}\qquad Q_{K}(P)=\frac{1}{N}\sum\limits_{i=1}^{N}\|\overline{T}_{K}^{(i)}-P\overline{T}_{K}^{(i)}\|_{L^{2}(\rho)}^{2}$
As remark in [43] it is well known that PCA amounts to solve
$P^{\leq q}\in\operatorname*{argmin}\limits_{P\in\mathcal{P}_{q}}Q(P)=\operatorname*{argmin}_{P\in\mathcal{P}_{q}}E(P),\quad\mbox{with}\quad E(P)=-\langle\Sigma,P\rangle_{HS},$
where $\langle\cdot,\cdot\rangle_{HS}$ is the Hilbert-Schmidt scalar product defined for linear operators $S,T:L^{2}(\rho)\rightarrow L^{2}(\rho)$ as $\langle S,T\rangle_{HS}=tr(S^{*}T)$. In the same way, we have that:
$P_{K}^{\leq q}\in\operatorname*{argmin}\limits_{P\in\mathcal{P}_{q}}E_{K}(P),\quad\mbox{ where }E_{K}(P)=-\langle\Sigma_{K},P\rangle_{HS}.$
We can finally define the excess risk of the PCA projector $P_{K}^{\leq q}$ with:
$\mathcal{E}^{PCA}_{q}=E(P_{K}^{\leq q})-E(P^{\leq q}).$
The following result allows to control the above excess risk with an upper bound depending on the quantization error $\varepsilon_{K}$.
Proposition 4.7.
Suppose that the $T^{(i)}$’s are $L$-Lipschitz for all $1\leq i\leq N$, then
$\mathcal{E}_{q}^{PCA}\leq 8R\sqrt{qL\ \mathrm{diam}(\mathcal{X})}\varepsilon_{K}^{1/4},$
with $R=\max\limits_{x\in\mathcal{X}}\|x\|$.
@@ 5 Numerical experiments
In our experiments, we distinguish the two quantization approaches as follows: $\tilde{K}$-LOT and $\tilde{K}$-KME refer to the LOT embedding and KME of $(\tilde{\nu}^{(i)}_{K})_{i=1}^{K}$ (3.1) obtained from the quantization of each $\mu^{(i)}$, while $\overline{K}$-LOT and $\overline{K}$-KME refer to the LOT embedding and KME of $(\overline{\nu}^{(i)}_{K})_{i=1}^{K}$ (3.4) obtained from the mean-measure quantization. Here, we aim to show that our methods enable fast computation of machine learning tasks while preserving the main information of the measures. We demonstrate their effectiveness by comparing our approaches with several methods:
  1. 1.
    KME with RFF. Random Fourier Features (RFF) [42] allow to efficiently approximate the KME by defining a feature map $\tilde{\varphi}(x)\in\mathbb{R}^{s}$ such that $\tilde{\varphi}(x)^{T}\tilde{\varphi}(y)\approx k(x,y)$. More precisely, this feature map is given by:
    $\tilde{\varphi}(x)=\bigl{(}\sin(\omega_{1}^{T}x),\cdots,\sin(\omega_{s/2}^{T}x),\cos(\omega_{1}^{T}x),\cdots,\cos(\omega_{s/2}^{T}x)\bigr{)},$
    where the $\omega_{i}$’s are independently sampled from a distribution related to the kernel $k$. In this framework, the KME therefore becomes the following mapping between probability measures and $\mathbb{R}^{s}$:
    $\tilde{\phi}(\mu)=\int_{\mathcal{X}}\tilde{\varphi}(x)\mathrm{d}\mu(x).$
    In order to fairly compare results, we choose $s=K$. In our experiments, we use the radial basis kernel (RBF) defined as:
    $k(x,y)=\exp\Bigl{(}-\frac{\|x-y\|^{2}}{2\sigma^{2}}\Bigr{)},$
    where $\sigma$ is a bandwidth parameter. The choice $\sigma$ is made via the median heuristic $\sigma^{2}=\mathrm{median}\{\|X^{(i)}_{j}-X^{(i)}_{l}\|^{2}\}.$
  2. 2.
    $K$-Nys-KME. The method proposed in [13], called Nyström Kernel Mean Embedding, proposes to estimate KME with a Nyström approximation using a small subset of size $K$ (echoing the parameter we deal with in quantization) of the data. More precisely, their approach consists in sampling $K$ points $(x_{k})_{1\leq k\leq K}$ from $\mu$ and find the optimal weights $(\alpha_{k})_{1\leq k\leq K}$ such that the estimator
    $\phi_{K-\mathrm{Nys}}(\mu)=\sum\limits_{k=1}^{K}\alpha_{k}k(x_{k},\cdot)$
    approximates at best the true KME $\phi(\mu)=\int_{\mathcal{X}}k(x,\cdot)\mathrm{d}\mu(x)$.
  3. 3.
    Random subset. Rather than selecting $K$ points with quantization, one could randomly sample $K$ points $(x^{(i)},\cdots,x^{(i)}_{K})$ from $\mu^{(i)}$ and construct the corresponding “non-optimal quantized measure” $\nu^{(i)}_{K}=\sum_{k=1}^{K}\mu^{(i)}(V_{x^{(i)}_{k}})\delta_{x^{(i)}_{k}}$ and then compute the LOT or KME embeddings of these discrete measures supported on a small number of points.
@@ 5.1 Computational cost
For a discrete measure, we solve the quantization problem (2.9) using Lloyd’s algorithm [33] and an initialization based on $K$-means+++++ + [4]. The time complexity of the Lloyd’s algorithm being linear in the number of data points [24], it follows that for discrete measures $\mu_{1},\ldots,\mu_{N}$ supported on $m_{1},\ldots,m_{N}$ points respectively, the computational cost for constructing the quantized measures $(\nu^{(i)}_{K})_{i=1}^{K}$ by either optimal quantization of each input measure or mean-measure quantization is $O(Kd\sum_{i=1}^{N}m_{i})$. Nevertheless, as the support of $\overline{\mu}$ is very large in applications, the mean measure quantization can be done on $\frac{1}{N}\sum_{i=1}^{N}m_{i}$ points randomly sampled from $\overline{\mu}$. Indeed, we have observed that subsampling the mean-measure does not deteriorate the performances of our experiments compared to using all points. Then, the computation cost of mean-measure quantization becomes $O(Kd\frac{1}{N}\sum_{i=1}^{N}m_{i})$ and is advantageous over the optimal quantization of each input measure.
In order to compute the LOT embedding, we solve the discrete OT problem (2.1) between $m_{0}$ samples $y_{1},\cdots,y_{m_{0}}$ of $\rho$ and $m_{i}$ samples $x^{(i)}_{1},\cdots,x^{(i)}_{m_{i}}$ of $\mu^{(i)}$ using a standard OT solver [40, 18]. Then, as the optimal map $T_{\rho}^{\mu^{(i)}}$ in (2.4) might not exist, it is classical [16] to compute an approximation through barycentric projection. Let $P^{(i)}\in\mathbb{R}^{m_{0}\times m_{i}}$ be an optimal transport plan, solution of (2.1), between discretized $\rho$ and $\mu^{(i)}$. Then, one defines the barycentric projection map as
$T^{(i)}:y_{j}\mapsto m_{0}\sum\limits_{l=1}^{m_{i}}P^{(i)}_{jl}x^{(i)}_{l}.$
The overall computational complexity of $\overline{K}$-LOT when using $m_{0}$ samples from the reference measure $\rho$ is thus $O(Kd\frac{1}{N}\sum_{i=1}^{N}m_{i})+O(N(K+M)Km_{0}\log(K+m_{0}))$ which is significantly smaller than the one of LOT from the raw input measures that scales as $O(\sum_{i=1}^{N}(m_{i}+m_{0})m_{i}m_{0}\log(m_{i}+m_{0}))$.
@@ 5.2 Synthetic dataset : shifts and scalings of a reference measure
$\mu^{(i)}=(\Sigma_{i}^{1/2}\mathrm{id}+b_{i})_{\#}\rho,$ (5.1)
where $\Sigma_{i}\in\mathbb{R}^{d\times d}$ is a positive semi-definite matrix and $b_{i}\in\mathbb{R}^{d}$. We choose $X\sim\rho$ such that $X=R\frac{Z}{\|Z\|}$, with $R\sim\mathrm{Unif}([0,1])$ and $Z\sim\mathcal{N}(0,I_{d})$. In that case, we have an explicit formulation of the pairwise inner-products induced by LOT embedding and KME, see Proposition C.1 in Appendix C. We can therefore exactly compute the so-called true matrix $G_{\mu}^{\phi}$. In order to compare this matrix to the one computed from our quantizations schemes of empirical measures, we numerically sample the distributions $\mu^{(i)}$’s in the following way : for each $1\leq i\leq N$, we first sample $m_{i}$ points $(x_{j})_{1\leq j\leq m_{i}}$ from the measure $\rho$ by sampling $z_{j}\sim\mathcal{N}(0,I_{d})$ and $r_{j}\sim\mathrm{Unif}([0,1])$ and computing $x_{j}=r_{j}\frac{z_{j}}{\|z_{j}\|}$. This allows to sample points from the unit ball in $\mathbb{R}^{d}$. Samples from $\mu^{(i)}$ are then obtained by the pushforward operation in (5.1).
We first compare in Figure 2 the computational costs of both quantization methods and observe that the mean-measure quantization approach is faster. Additionally, we depict in Figure 4 the evolution of $\|G^{\phi}_{\mu}-G^{\phi}_{\nu_{K}}\|_{F}$ for different values of $K$ and $\phi$ denoting either the LOT or the KME embedding, whose convergence was shown in Proposition 4.6. As expected (see Remark 3.6), the quantization of each measure method yields better approximations than mean-measure quantization. In Figure 3 (resp. Figure 5 in Appendix C), we visualize the projections on the first two components of PCA for $K$-LOT (resp. $K$-KME) of both quantization methods and compare them to the PCA computed from the true Gram matrices $G_{\mu}^{\phi}$ computed in Proposition C.1. We observe that even with a small value $K=32$, the PCA visualizations on quantized embedded measures mimic the ones on raw embedded input measures.
@@
Figure 3: Synthetic dataset on shifts and scalings. Projection of the data (colored dots) onto the first two components of PCA after $\overline{K}$-LOT (top) and $\tilde{K}$-LOT (bottom) and comparison to the LOT PCA (right) computed from the true Gram matrix (see Prop.C.1).
Figure 4: Synthetic dataset on shifts and scalings. Approximation error of the Gram matrices for $\tilde{K}$-LOT, $\overline{K}$-LOT, $\tilde{K}$-KME, $\overline{K}$-KME in function of $K$.
@@ 5.3 Flow cytometry dataset
Table 1: Flow cytometry dataset. LDA classification accuracies and execution times after 10-component PCA on the methods with $K=64$.
Method Accuracy (Lab) Accuracy (Type) Time (s)
$\overline{K}$-LOT 100 94 30
$\tilde{K}$-LOT 100 81 103
Random subset of size $K$ + LOT 100 77 25
Random subset of size $K$ + KME 100 69 101
$\overline{K}$-KME 100 85 105
$\tilde{K}$-KME 100 69 358
KME with RFF 83 52 5035
$K$-Nys-KME 100 73 36
@@
In this section, we use flow cytometry datasets provided in [48] and publicly available in Mendeley Data to illustrate the suitability of our method through a classification task. We have $N=108$ cytometry measures (or point cloud) which come from two different health care centers : Marburg and Dresden. In Marburg, the data consists of diagnostic samples of peripheral blood (pB), healthy bone marrow (BM), or leukemic bone marrow. The Dresden dataset consists of diagnostic samples of peripheral blood and healthy bone marrow. Two types of labels can be distinguished: data are differentiated either by the healthcare centers from which they were analyzed, or by their types (e.g., peripheral blood, healthy bone marrow, or leukemic bone marrow). Each measure contains from 100,000 to 1,000,000 points in dimension $d=10$, which prevents the use of classical OT, that is, without a quantization step. For $K$-LOT, we sample $m_{0}=1000$ points from the reference measure chosen as the uniform measure on $[0,1]^{d}$ and for $K$-KME, we use the Gaussian kernel with bandwidth parameter $\sigma=1$. The embedded measures live in a high-dimensional Hilbert space (e.g. with $K$-LOT and $K=128$, the ambient space is $\mathbb{R}^{1280}$). In order to keep the most relevant information, we perform a 10-components PCA on the embedded data with respect to the proper Hilbert space $\mathcal{H}$. Then, we train two classifiers (one for each type of label) on 75% of the data and test the results on the remaining data with LDA. Accuracy scores are displayed in Table 1 for $K=64$. Additional experiments for different values of $K$ can be found in Appendix C.3. The quantization methods achieve in particular perfect accuracy scores for predicting the health care center (denoted as LAB) in only a few minutes. We observe that $K$-LOT consistently outperforms $K$-KME. More generally, both approaches outperform the other benchmarked methods. KME with RFF notably performs poorly. These results also underscore the relevance of choosing meaningful centroids with quantization rather than relying on random sampling. Comparing the quantization methods, we observe that the classifier performs better on mean-measure quantization than on the quantization of each measure whereas the latter is roughly ten times slower than the former.
@@
Additionally, we visualize the projections of the data on the first components of PCA in Figures 6 and 7 of the Appendix C. In these figures, it is clear that the representations stabilize when $K\geq 32$. This shows that our $K$-quantization step gives good results even for small values of $K$. We also observe that the mean-measure quantization and the quantization of each measure yield similar results.